# Multimodal Survival Modeling in the Age of Foundation Models

### v2

@@ RESULTS
While unimodal models were able to achieve a peak C-index of approximately 0.75, we find that all multimodal fusion models are able to surpass unimodal results at their corresponding PCA reductions (Tables 3 and 4). We again observe a plateau of survival model performance at PCA = 256 and thus report results using this dimensionality reduction for embeddings. Importantly, each modality is independently modeled at a given PCA dimensionality before the predicted unimodal risk scores are subsequently used as input features to the multimodal fusion model (see Methods for more details).
@@
Using this simple, late multimodal fusion technique, we find that fusion of expression with histology data, expression with report data, and histology with report data result in mean cross-validated C-index = 0.774, 0.778, and 0.780, respectively (Table 3). Fusion of all three of expression, histology, and reports results in the best embedding-based model performance of C-index = 0.788. Survival modeling with these embedding-based modalities are not only additive, but also can be further enhanced with clinical features, such as demographics, achieving our greatest C-index = 0.795 (Table 4).
When examining multimodal model performance subset by cancer type in Table 2, we find that multimodal fusion consistently performs better than any other modality within a given cancer type. This is in contrast to the unimodal models where the best unimodal model differs depending on the cancer. Furthermore, across the 8 most common cancer types in our data, we find that the multimodal model’s worst performance (C-index = 0.626 for LUSC) does not cross below a C-index of 0.6, whereas each unimodal model does.
Pathology report summarization focuses survival information
Modality Cancer Type (TCGA Project)
@@
BRCA KIRC UCEC THCA LGG HNSC LUSC LUAD
Original 0.612 0.645 0.575 0.593 0.630 0.574 0.540 0.593
Summarized 0.641 0.698 0.723 0.619 0.716 0.582 0.556 0.645
Multimodal 0.720 0.752 0.758 0.787 0.840 0.650 0.626 0.687
@@
Table 5: Summarization of pathology reports improves pan-cancer survival prediction. Averaged C-index from 5-fold cross-validation for the top 8 most prevalent cancer types in TCGA. All text embedded using BioMistral. Summarized text generated using Llama-3.1-8B-Instruct. Multimodal model incorporates unimodal summarized text model. Results reported using PCA to 256 dimensions for all foundation model derived embeddings.

### v3

@@ 3.3 Multimodal fusion improves cancer modeling and survival prediction
While unimodal models were able to achieve a peak C-index of approximately 0.75, we find that all multimodal fusion models are able to surpass unimodal results at their corresponding PCA reductions (Tables LABEL:tab:best-multimodal and LABEL:tab:fm-multimodal). We again observe a plateau of survival model performance at PCA = 256 and thus report results using this dimensionality reduction for embeddings. Importantly, each modality is independently modeled at a given PCA dimensionality before the predicted unimodal risk scores are subsequently used as input features to the multimodal fusion model (Section 2.4).
@@
Using this simple, late multimodal fusion technique, we find that fusion of expression with histology data, expression with report data, and histology with report data result in mean cross-validated C-index = 0.774, 0.778, and 0.780, respectively (Table LABEL:tab:fm-multimodal). Fusion of all three of expression, histology, and reports results in the best embedding-based performance of C-index = 0.788. Survival modeling with these embedding-based modalities are not only additive, but also can be further enhanced with clinical features, such as demographics, achieving our greatest C-index = 0.795 (Table LABEL:tab:best-multimodal). The additive effect of multimodal fusion is further observed in both mean AUCC,D and IBS (Tables LABEL:tab:time-dependent-auroc and LABEL:tab:integrated-brier-score, respectively). Importantly, these modalities are encoding information beyond simply cancer type, as any combination of modalities with cancer type improves upon the results of using cancer type alone (Table LABEL:tab:combinations-w-cancer-type).
[…]
@@
tab:per-project-8 Modality Cancer Type (TCGA Project) BRCA KIRC UCEC THCA LGG HNSC LUSC LUAD Demographics 0.637 0.600 0.593 0.885 0.727 0.525 0.542 0.530 RNA-seq 0.573 0.673 0.643 0.619 0.790 0.560 0.579 0.620 Histology 0.648 0.685 0.703 0.749 0.753 0.636 0.598 0.617 Text 0.641 0.698 0.723 0.619 0.716 0.582 0.556 0.645 Multimodal 0.720 0.752 0.758 0.787 0.840 0.650 0.626 0.687
@@
Table 3: Multimodal fusion improves survival prediction within the 8 most prevalent cancer types in the TCGA. Average cross-validated C-index of models using PCA to 256 dimensions for all foundation model derived embeddings. Multimodal model used fusion of all unimodal modalities.
Additionally, as before, the ability to model survival within each cancer type further substantiates this claim, as the unimodal model based on cancer type cannot predict survival within a single cancer type. When examining multimodal model performance subset by cancer type in Tables LABEL:tab:per-project-8 and LABEL:tab:per-project, we find that multimodal fusion consistently performs better than any other modality within a given cancer type for the most prevalent cancer types (see Appendix C for additional details on evaluations within cancer types). This is in contrast to the unimodal models where the best unimodal model differs depending on the cancer.
Interpretability. Under our framework for multimodal fusion, a natural interpretation of the multimodal model arises. As the input to the multimodal CoxPH model are z-score normalized unimodal risk scores, the hazard ratios of the trained multimodal model can be interpreted relative to the standard deviations of the unimodal predicted risks. In other words, the hazard ratio for a given modality is the relative risk of death for a one standard deviation increase in that modality’s unimodal predicted risk. Table LABEL:tab:hazard-ratios presents the hazard ratios of the 5-modality model (using unimodal models over PCA = 256 for high-dimensional embeddings). We find that the high-dimensional inputs confer greater relative risk, recapitulating our observations that these modalities are informative beyond tabular features.
3.4 Pathology report summarization focuses cancer information
One of our key findings is that LLM summarization of pathology reports drastically improves survival prediction. Specifically, we zero-shot prompt Llama-3.1-8B-Instruct to summarize pathology reports with a focus on microscopic descriptions, test results, diagnoses, and clinical history. See Section 2.2 for more details on our summarization method.
Using Llama generated summaries, we embed the summarized reports using BioMistral. As compared with BioMistral embeddings of the original, unsummarized reports, we find that the summarized reports are able to better predict survival (C-index = 0.752 for summarized vs C-index = 0.694 for unsummarized). Using the predicted risk scores, we stratify the study cohort into low and high risk groups and observe that summarized reports result in enhanced risk stratification (Figure LABEL:fig:summarization).
Not only does summarization reduce the token length of the original report, it also corrects typographical errors that may result from text extraction from the source scanned PDFs of these reports. In addition, prompting the summarization to focus on certain aspects of the report may better extract information most relevant to survival prediction. With these effects in combination, we observe that embeddings of summarization outperforms embeddings of the original report across the 8 most prevalent cancer types (Table LABEL:tab:per-project-text).
3.5 Hallucination correction does not impact cancer risk stratification
While we empirically find that summarization improves survival prediction, hallucinations are an important consideration when using LLM generated text. To test whether factually incorrect information in the generated summaries impact downstream survival modeling, we manually correct generated summaries for these hallucinations.
To that end, we develop and share a lightweight tool to facilitate comparison of original and summarized reports. A screenshot of the tool is presented in Figure LABEL:fig:comparison-tool. Using this utility, we review 40 randomly sampled reports contained within a single cross-validation fold. We ultimately apply minor corrections to 24 of 40 reports. See Appendix A for additional details, such as methodology and categorization of manual corrections.
After manual correction, we re-embed the corrected summaries and do survival prediction using the same CoxPH model trained on uncorrected summarized embeddings. For the subset of corrected reports, we compare against the corresponding uncorrected summaries (Figure LABEL:fig:hallucination). We empirically find that risk stratification of the sampled subset does not change based on correction of hallucinations. This may be because the corrections on average are qualitatively minor compared to the majority of the summary (Table LABEL:tab:correction-categories). We hypothesize that the salient information for predicting cancer survival, such as cancer severity, extent of invasion or metastasis, remained unchanged, though further work is needed.
3.6 Impact of domain specificity on cancer modeling and survival prediction
Our final set of experiments were ablations to test the impact of foundation model domain specificity for downstream embedding-based modeling. For space, extended results are presented in Appendix D. In brief, domain adaptation is particularly important for data modality specificity (BulkRNABert is notably better than UCE for RNA-seq data, Figure LABEL:fig:expr-models), however specificity to biomedical text (BioMistral vs Mistral) matters less than summarization (Figure LABEL:fig:text-models).
[…]
fig:summarization Refer to caption